# Statistical Learning of Distributionally Robust Stochastic Control in Continuous State Spaces

### v1

@@ Abstract
We explore the control of stochastic systems with potentially continuous state and action spaces, characterized by the state dynamics $X_{t+1}=f(X_{t},A_{t},W_{t})$. Here, $X$, $A$, and $W$ represent the state, action, and exogenous random noise processes, respectively, with $f$ denoting a known function that describes state transitions. Traditionally, the noise process $\left\{{W_{t},t\geq 0}\right\}$ is assumed to be independent and identically distributed, with a distribution that is either fully known or can be consistently estimated. However, the occurrence of distributional shifts, typical in engineering settings, necessitates the consideration of the robustness of the policy. This paper introduces a distributionally robust stochastic control paradigm that accommodates possibly adaptive adversarial perturbation to the noise distribution within a prescribed ambiguity set. We examine two adversary models: current-action-aware and current-action-unaware, leading to different dynamic programming equations. Furthermore, we characterize the optimal finite sample minimax rates for achieving uniform learning of the robust value function across continuum states under both adversary types, considering ambiguity sets defined by $f_{k}$-divergence and Wasserstein distance. Finally, we demonstrate the applicability of our framework across various real-world settings.
@@ 1 Introduction
Stochastic control formulations are extensively utilized in the modeling, design, and optimization of systems influenced by probabilistic dynamics. These formulations play a crucial role across various fields within operations research and management disciplines. Notable applications of stochastic control can be seen in finance (Merton, 1976), communication systems (Yüksel and Başar, 2013), manufacturing and operations management (Tse and Grossglauser, 1997; Porteus, 2002), as well as energy systems (Foschini and Miljanic, 1993). A key aspect common to these applications is the use of a continuous state space, which provides a robust and flexible environment for formulating complex system dynamics. This facilitates the development of realistic dynamic models that accurately reflect the underlying systems, thereby enhancing system management and operational efficiency.
The underlying state dynamics of a large class of stochastic control problems can be described by the following recursion

### v2

# Learning Optimal Distributionally Robust Stochastic Control in Continuous State Spaces

@@ Abstract
We study data-driven learning of robust stochastic control for infinite-horizon systems with potentially continuous state and action spaces. In many managerial settings–supply chains, finance, manufacturing, services, and dynamic games–the state-transition mechanism is determined by system design, while available data capture the distributional properties of the stochastic inputs from the environment. For modeling and computational tractability, a decision maker often adopts a Markov control model with i.i.d. environment inputs, which can render learned policies fragile to internal dependence or external perturbations. We introduce a distributionally robust stochastic control paradigm that promotes policy reliability by introducing adaptive adversarial perturbations to the environment input, while preserving the modeling, statistical, and computational tractability of the Markovian formulation. From a modeling perspective, we examine two adversary models–current-action-aware and current-action-unaware–leading to distinct dynamic behaviors and robust optimal policies. From a statistical learning perspective, we characterize optimal finite-sample minimax rates for uniform learning of the robust value function across a continuum of states under ambiguity sets defined by the $f_{k}$-divergence and Wasserstein distance. To efficiently compute the optimal robust policies, we further propose algorithms inspired by deep reinforcement learning methodologies. Finally, we demonstrate the applicability of the framework to real managerial problems.
@@ 1 Introduction
Stochastic optimal control formulations are extensively utilized in the modeling, design, and optimization of systems influenced by probabilistic dynamics. These formulations play a crucial role across various fields within operations research and management disciplines. Notable applications of stochastic control can be seen in finance [26], supply chain management [44], communication systems [55], manufacturing and operations management [42, 33], as well as energy systems [11]. Modeling such systems naturally leads to the use of continuous state spaces, supporting high-resolution dynamic models that accurately capture the underlying system dynamics.
The state transition dynamics of a broad class of stochastic control problems can be expressed through the following recursive relation